OpenStudy (anonymous):

Solve this system of equations using the addition method x-y=6 and -3y=x+22

OpenStudy (goformit100):

First of all @allmightystew : A Warm Welcome to 'Open Study'. Please Read CoC (compulsory to read by all "Open Study" users) : http://openstudy.com/code-of-conduct

OpenStudy (radar):

x = y + 6 and x =-3y - 22 x = x y + 6 = -3y -22 4y = -22 - 6 4y = -28 y = -7 x=-7 + 6 = -1 I don't believe this method is what they wanted. (But it works and is the simplest)

OpenStudy (anonymous):

thank you

OpenStudy (radar):

Here is the addition method. x - y = 6 now we will add the 2nd equation to the first equation -x-3y = 22 adding we get ------------ -4y = 28 y = -7 now substitute the -7 value for y in the first equation getting x - (-7) = 6 x + 7 = 6 x = 6 - 7 x = -1 Same answers as the first go around.

OpenStudy (radar):

Did you follow with understanding @allmightystew ?

OpenStudy (anonymous):

yup thank you i needed a refresh on this and this helped alot :)

OpenStudy (radar):

You're welcome and come back any time.